# ReSearch: A Multi-Stage Machine Learning Framework for Earth Science Data Discovery

### v1

@@ Abstract
The rapid expansion of Earth Science data from satellite observations, reanalysis products, and numerical simulations has created a critical bottleneck in scientific discovery, namely identifying relevant datasets for a given research objective. Existing discovery systems are primarily retrieval-centric and struggle to bridge the gap between high-level scientific intent and heterogeneous metadata at scale. We introduce ReSearch, a multi-stage, reasoning-enhanced search framework that formulates Earth Science data discovery as an iterative process of intent interpretation, high-recall retrieval, and context-aware ranking. ReSearch integrates lexical search, semantic embeddings, abbreviation expansion, and large language model reranking within a unified architecture that explicitly separates recall and precision objectives. To enable realistic evaluation, we construct a literature-grounded benchmark by aligning natural language intent with datasets cited in peer-reviewed Earth Science studies. Experiments demonstrate that ReSearch consistently improves recall and ranking performance over baseline methods, particularly for task-based queries expressing abstract scientific goals. These results underscore the importance of intent-aware, multi-stage search as a foundational capability for reproducible and scalable Earth Science research.
@@ 1 Introduction
Recent advances in machine learning offer new opportunities to address these limitations. Semantic embedding models and large language models (LLMs) have demonstrated strong capabilities in representing meaning, handling linguistic variability, and interpreting natural language queries. These tools provide a promising foundation for bridging the gap between scientific intent and technical metadata. However, purely neural approaches often lack transparency, stability, and explicit control over recall and relevance. These limitations highlight the need for discovery frameworks that combine machine learning with principled search strategies, explicitly encoding the stages of reasoning that underlie scientific inquiry.
@@ 2 Related Work
A growing body of work has explored the use of machine learning and semantic technologies to enhance Earth Science data discovery and interoperability [1, 3]. Major data portals and federated archives provide centralized access to observational and model datasets, while recent approaches have investigated semantic representations, embeddings, and knowledge graphs to bridge heterogeneity across repositories. Although these systems improve data accessibility, they are primarily designed for human-driven retrieval and offer limited support for intent-aware or reasoning-oriented discovery.
@@
Beyond knowledge-graph-based systems, embedding-based semantic retrieval has been employed to improve coverage across large, lexically diverse archives. For example, Qiu et al. [4] proposed an ontology-enhanced embedding model to support keyphrase extraction from geoscience literature. Ramachandran et al. [5] incorporated prediction-based embeddings to augment metadata in NASA Earth science repositories. More recently, Weckmüller et al. [6] designed a multilingual embedding-based search framework for geo-textual datasets. These approaches enhance recall and language robustness, but often conflate retrieval and ranking, and offer limited mechanisms for integrating structured constraints or domain-specific reasoning.
Recent trends increasingly emphasize modeling data discovery as a multi-stage reasoning process, where components such as intent interpretation, candidate generation, and contextual ranking are decoupled. Such designs offer greater flexibility across heterogeneous repositories and can integrate with both semantic and agent-based systems. This work builds on these directions, seeking to extend the reasoning capacity of discovery systems beyond rigid taxonomies and single-step retrieval.
@@ 4.1 Data Sources
To establish a robust foundation for Earth Science research, we have integrated four primary data repositories. These include the NASA Common Metadata Repository (CMR), NOAA OneStop, the Coupled Model Intercomparison Project Phase 6 (CMIP6), and the ERA5 (ECMWF Reanalysis v5) dataset. These sources collectively represent a vast and diverse collection of Earth observation and simulation data, ranging from satellite measurements to numerical model outputs.
@@ 4.2 Search Strategy Design
The final stage refines the candidate set using an LLM-based reranker. By evaluating the detailed metadata of retrieved datasets (including titles and summaries) against the nuanced context of the user’s original query, this stage assigns a relevance score to each dataset. This process effectively promotes the most pertinent resources to the top of the results, filtering out noise introduced during the high-recall retrieval phase.
@@ 4.3 Evaluation Dataset Construction
Table 2: Evaluation Dataset Statistics
Item Count
Sampled academic papers 15
Extracted keywords 80
Task-based queries (“I want to…”) 30
Datasets mentioned 49
@@ 5.2 Main Results
AutoClimDS, which relies on knowledge graph traversal from auxiliary concept nodes, achieves consistently low recall across all cutoff thresholds ($\textnormal{R@100}\leq 0.02$). This suggests that the bipartite graph structure, while interpretable, cannot recover datasets that fall outside predefined semantic categories. The relatively higher MRR (0.071) indicates that when AutoClimDS does retrieve a relevant dataset, it tends to appear early in the ranked list; however, the overall coverage remains insufficient for practical discovery scenarios.
BM25 achieves strong recall on keyword queries.
For keyword-based queries, BM25 achieves the highest recall at all cutoff levels ($\textnormal{R@100}=0.32$), outperforming embedding-based retrieval ($\textnormal{R@100}=0.27$). This result reflects the nature of keyword queries, which contain precise technical terms that match well with lexical retrieval. BM25’s effectiveness in this setting underscores the continued importance of exact term matching for queries where users already know the relevant vocabulary.
Embedding retrieval improves ranking quality.
Although embedding retrieval achieves lower recall than BM25 on keyword queries, it demonstrates higher MRR (0.065 vs. 0.040), indicating that semantically relevant datasets are ranked higher when they are retrieved. This suggests that embedding-based methods capture semantic similarity beyond lexical overlap, which benefits ranking even when overall coverage is lower. For task-based queries, embedding retrieval shows comparable or slightly better performance than BM25, reflecting its ability to handle the vocabulary mismatch between abstract research objectives and technical metadata.
Task-based queries remain challenging.
Both BM25 and embedding retrieval exhibit lower recall on task-based queries compared to keyword queries. This performance gap highlights the semantic gap between high-level scientific intent and low-level metadata descriptors. Task-based queries such as “I want to study the impact of Arctic sea ice loss on mid-latitude weather” require reasoning over implicit data requirements, which neither lexical nor semantic retrieval can fully address without additional query understanding or reranking stages.
@@ 5.3 Ablation Study: Effect of Abbreviation Expansion
At R@100, the improvement from abbreviation expansion is more modest (BM25: $0.31\rightarrow 0.32$; Embedding: $0.25\rightarrow 0.27$). This pattern suggests that abbreviation expansion primarily helps surface relevant datasets earlier in the ranked list, rather than recovering entirely new relevant datasets. The practical implication is that abbreviation expansion is particularly valuable for user-facing applications where users examine only the top results.
@@ 6 Conclusion
The proposed framework integrates complementary techniques from information retrieval and machine learning, including lexical matching, semantic embedding search, and abbreviation expansion, with a modular architecture designed to support large language model reasoning in downstream ranking stages. This hybrid design improves robustness to metadata heterogeneity and linguistic variation, which are pervasive challenges in Earth Science data ecosystems. Through a literature-grounded evaluation derived from peer-reviewed studies, we demonstrated that the retrieval components of ReSearch outperform baseline methods, particularly for task-based queries that reflect abstract scientific objectives rather than dataset-specific terminology.
@@ 3. What spatial/temporal aspects matter?
B.3 Paper Information Extraction Prompt

### v2

# arXiv:2601.14176v2[cs.DB] 02 Mar 2026

@@ Abstract
The rapid expansion of Earth Science data from satellite observations, reanalysis products, and numerical simulations has created a critical bottleneck in scientific discovery, namely identifying relevant datasets for a given research objective. Existing discovery systems are primarily retrieval-centric and struggle to bridge the gap between high-level scientific intent and heterogeneous metadata at scale. We introduce ReSearch, a multi-stage, reasoning-enhanced search framework that formulates Earth Science data discovery as an iterative process of intent interpretation, high-recall retrieval, and context-aware ranking. ReSearch integrates lexical search, semantic embeddings, abbreviation expansion, and large language model reranking within a unified architecture that explicitly separates recall and precision objectives. To enable realistic evaluation, we construct a literature-grounded benchmark by aligning natural language intent with datasets cited in peer-reviewed Earth Science studies. Experiments demonstrate that ReSearch consistently improves recall and ranking performance over baseline methods, particularly for task-based queries expressing abstract scientific goals. These results demonstrate the importance of intent-aware, multi-stage search as a foundational capability for reproducible and scalable Earth Science research.
@@ 1 Introduction
Recent advances in machine learning offer new opportunities to address these limitations. Semantic embedding models and large language models (LLMs) have demonstrated strong capabilities in representing meaning, handling linguistic variability, and interpreting natural language queries. These tools provide a promising foundation for bridging the gap between scientific intent and technical metadata. However, purely neural approaches often lack transparency, stability, and explicit control over recall and relevance. These limitations indicate the need for discovery frameworks that combine machine learning with principled search strategies, explicitly encoding the stages of reasoning that underlie scientific inquiry.
@@
Figure 1: Overview of the ReSearch framework. User queries enter the search engine from a web page or GPT-based agents. Inside the engine, keyword-based and embedding-based recall run in parallel, followed by LLM reranking. Results are returned for direct use or further processing by downstream agents.
@@ 2 Related Work
A growing body of work has explored the use of machine learning and semantic technologies to improve Earth Science data discovery and interoperability [1, 4]. Major data portals and federated archives provide centralized access to observational and model datasets, while recent approaches have investigated semantic representations, embeddings, and knowledge graphs to bridge heterogeneity across repositories. Although these systems improve data accessibility, they are primarily designed for human-driven retrieval and offer limited support for intent-aware or reasoning-oriented discovery.
@@
Beyond knowledge-graph-based systems, embedding-based semantic retrieval has been employed to improve coverage across large, lexically diverse archives. For example, Qiu et al. [5] proposed an ontology-enhanced embedding model to support keyphrase extraction from geoscience literature. Ramachandran et al. [6] incorporated prediction-based embeddings to augment metadata in NASA Earth science repositories. More recently, Weckmüller et al. [7] designed a multilingual embedding-based search framework for geo-textual datasets. These approaches improve recall and language robustness, but often conflate retrieval and ranking, and offer limited mechanisms for integrating structured constraints or domain-specific reasoning.
Recent trends increasingly treat data discovery as a multi-stage reasoning process, where components such as intent interpretation, candidate generation, and contextual ranking are decoupled. Such designs offer greater flexibility across heterogeneous repositories and can integrate with both semantic and agent-based systems. This work builds on these directions, seeking to extend the reasoning capacity of discovery systems beyond rigid taxonomies and single-step retrieval.
@@ 4.1 Data Sources
To establish a solid basis for Earth Science research, we have integrated four primary data repositories. These include the NASA Common Metadata Repository (CMR), NOAA OneStop, the Coupled Model Intercomparison Project Phase 6 (CMIP6), and the ERA5 (ECMWF Reanalysis v5) dataset. These sources are a large and diverse collection of Earth observation and simulation data, ranging from satellite measurements to numerical model outputs.
@@ 4.2 Search Strategy Design
The final stage refines the candidate set using an LLM-based reranker (GPT-4o). The top $K_{c}$ ($K_{c}=50$) merged candidates from Stage 1 are formatted as a numbered list with truncated titles and summaries, then presented to the LLM with the user’s query. The LLM returns the indices of the $K_{r}$ ($K_{r}=10$) most relevant datasets in ranked order. The final result is constructed by placing these $K_{r}$ datasets first, followed by the remaining candidates in their original order, so that recall at higher cutoffs is preserved while early precision is maximized.
@@ 4.3 Evaluation Dataset Construction
To evaluate generalization at larger scale, we constructed a second benchmark from the NASA Earth Observation Knowledge Graph [3], which contains 8,058 satellite dataset nodes and 138,704 publication nodes, of which 19,871 publications reference at least one dataset. We ranked eligible publications by citation count, selected the top 1,000, and sampled every 10th paper to obtain 100 high-impact publications covering 206 dataset references. Since the Knowledge Graph provides only titles and abstracts, keywords and “I want to” statements were generated by GPT-4o, consistent with the first benchmark.
Table 2: Evaluation Dataset Statistics. NASA EO-KG refers to the NASA Earth Observation Knowledge Graph [3].
Item Literature NASA EO-KG
Sampled papers 15 100
Keyword-based queries 14 100
Task-based queries 28 200
Dataset references 49 206
@@ 5.2 Main Results
Rerank 0.13 0.15 0.19 0.29 0.094 0.047
NASA EO-KG Benchmark – Keyword-based Queries (n=100)
AutoClimDS 0.00 0.00 0.00 0.00 0.000 0.000
BM25 0.08 0.12 0.18 0.23 0.057 0.049
Embedding 0.09 0.14 0.18 0.24 0.057 0.047
Rerank 0.13 0.16 0.20 0.24 0.086 0.076
NASA EO-KG Benchmark – Task-based Queries (n=200)
AutoClimDS 0.00 0.00 0.00 0.00 0.000 0.000
BM25 0.09 0.10 0.15 0.19 0.068 0.054
Embedding 0.09 0.14 0.18 0.20 0.061 0.051
Rerank 0.13 0.16 0.17 0.21 0.081 0.063
@@
AutoClimDS, which relies on knowledge graph traversal from auxiliary concept nodes, achieves extremely low recall across all cutoff thresholds ($\textnormal{R@100}\leq 0.02$). Although the relatively higher MRR (0.071) indicates that when AutoClimDS does retrieve a relevant dataset it tends to appear early in the ranked list, the overall coverage is severely insufficient for practical discovery scenarios.
ReSearch outperforms baselines, followed by BM25 and Embedding.
For keyword-based queries, ReSearch achieves the best performance across all metrics ($\textnormal{R@100}=0.38$, $\mathrm{MRR}=0.172$, $\mathrm{MAP}=0.121$). BM25 ranks second ($\textnormal{R@100}=0.32$), while Embedding ranks third ($\textnormal{R@100}=0.27$). Notably, BM25 achieves higher recall than Embedding, indicating that exact lexical matching remains important for keyword queries; however, Embedding achieves higher MRR ($0.065$ vs. $0.040$), suggesting that semantic matching provides better ranking quality. For task-based queries, the same ranking holds: ReSearch leads ($\textnormal{R@100}=0.29$, $\mathrm{MRR}=0.094$), followed by BM25 ($\textnormal{R@100}=0.28$) and Embedding ($\textnormal{R@100}=0.23$). The performance gap between methods is smaller than in keyword queries, but the ranking order remains consistent.
Task-based queries are more challenging.
All methods exhibit lower performance on task-based queries compared to keyword-based queries. ReSearch’s R@100 drops from $0.38$ to $0.29$, BM25 from $0.32$ to $0.28$, and Embedding from $0.27$ to $0.23$. Task-based queries express high-level research objectives (e.g., “I want to study the impact of drought on agriculture”), which creates a semantic gap with dataset metadata. Bridging this gap is the core challenge that our multi-stage framework addresses.
5.3 Hybrid retrieval improves recall and ranking
To quantify the benefit of combining BM25 and embedding retrieval, we compare retrieval performance across three configurations: BM25 only, Embedding only, and BM25+Embedding fusion. Table 4 presents the results.
Table 4: Effect of Retriever Fusion (without reranking). Combining BM25 and Embedding retrieval improves recall at higher cutoffs.
Method R@10 R@20 R@50 R@100 MRR MAP
Keyword-based Queries (n=14)
BM25 0.12 0.15 0.27 0.32 0.040 0.033
Embedding 0.06 0.09 0.26 0.27 0.065 0.029
BM25+Embedding 0.05 0.18 0.33 0.38 0.098 0.043
Task-based Queries (n=28)
BM25 0.04 0.11 0.22 0.28 0.062 0.026
Embedding 0.07 0.12 0.19 0.23 0.049 0.031
BM25+Embedding 0.06 0.15 0.19 0.29 0.094 0.047
Combining BM25 and embedding retrieval yields substantial improvements in both recall and ranking quality. For keyword-based queries, hybrid retrieval achieves $\textnormal{R@100}=0.38$ compared to $0.32$ for BM25 alone and $0.27$ for embedding alone. MRR improves from $0.040$ (BM25) and $0.065$ (Embedding) to $0.098$, indicating that relevant datasets are ranked higher when both retrieval signals are combined. Similar patterns hold for task-based queries, where hybrid retrieval achieves the highest MRR ($0.094$) and the best R@100 ($0.29$ vs. $0.28$ for BM25 and $0.23$ for Embedding). The improvement demonstrates that BM25 and embedding-based methods capture complementary aspects of relevance: BM25 excels at exact lexical matching, while embeddings capture semantic similarity. By prioritizing datasets retrieved by both methods and interleaving the remaining results, the hybrid approach benefits from both signals.
5.4 Ablation Study: Effect of LLM Reranking
To quantify the benefit of LLM-based reranking, we fix the retriever to BM25+Embedding and compare performance with and without reranking. Table 5 presents the results.
Table 5: Effect of LLM Reranking (with BM25+Embedding retrieval). Reranking improves early precision without affecting recall.
Method R@10 R@20 R@50 R@100 MRR MAP
Keyword-based Queries (n=14)
Without Reranking 0.05 0.18 0.33 0.38 0.098 0.043
With Reranking 0.18 0.28 0.33 0.38 0.172 0.121
Task-based Queries (n=28)
Without Reranking 0.06 0.15 0.19 0.29 0.094 0.047
With Reranking 0.13 0.15 0.19 0.29 0.094 0.047
For keyword-based queries, reranking yields dramatic improvements: R@10 increases from $0.05$ to $0.18$, R@20 from $0.18$ to $0.28$, MRR from $0.098$ to $0.172$, and MAP from $0.043$ to $0.121$. This demonstrates that the LLM can effectively identify the most relevant datasets from the candidate pool and promote them to the top of the ranked list. Since users typically examine only the top 10–20 results, this improvement in early precision is critical for user experience. Meanwhile, R@50 and R@100 remain unchanged ($0.33$ and $0.38$, respectively), which is expected since reranking only reorders the top-50 candidates without introducing new datasets.
For task-based queries, R@10 improves from $0.06$ to $0.13$, but R@20, MRR, and MAP remain unchanged. This suggests that for queries expressing high-level research objectives, even LLM reasoning cannot fully bridge the semantic gap between query intent and dataset metadata. Nevertheless, the improvement in R@10 indicates that reranking still provides some benefit in surfacing the most relevant results at the very top of the list.
@@ 5.5 Ablation Study: Effect of Abbreviation Expansion
At R@100, the improvement from abbreviation expansion is more modest (BM25: $0.31\rightarrow 0.32$; Embedding: $0.25\rightarrow 0.27$). This pattern suggests that abbreviation expansion primarily helps surface relevant datasets earlier in the ranked list, rather than recovering entirely new relevant datasets. The practical implication is that abbreviation expansion is particularly useful for user-facing applications where users examine only the top results.
@@ 6 Conclusion
The proposed framework integrates complementary techniques from information retrieval and machine learning, including lexical matching, semantic embedding search, and abbreviation expansion, with a modular architecture designed to support large language model reasoning in downstream ranking stages. This hybrid design improves tolerance of metadata heterogeneity and linguistic variation, which are pervasive challenges in Earth Science data ecosystems. Through a literature-grounded evaluation derived from peer-reviewed studies, we demonstrated that the retrieval components of ReSearch outperform baseline methods, particularly for task-based queries that reflect abstract scientific objectives rather than dataset-specific terminology.
@@
Acknowledgments
HY was partially supported by the US National Science Foundation under awards IIS-2520978, GEO/RISE-5239902, the Office of Naval Research Award N00014-23-1-2007, DOE (ASCR) Award DE-SC0026052, and the DARPA D24AP00325-00. Approved for public release; distribution is unlimited.
@@ 3. What spatial/temporal aspects matter?
B.3 Reranking Prompt
This prompt instructs the LLM to rerank candidate datasets by relevance to the user’s query.
You are an expert Earth science dataset retrieval reranker.
Given a user query and numbered candidate datasets, rank them by relevance.
Rules:
- Prefer datasets that directly measure the key variables/phenomena
  implied by the query.
- Only return numbers from the given candidates. Do NOT invent new numbers.
- Return the TOP {top_n} most relevant candidates only.
Return JSON: {"ranked": [3, 1, 5, ...]}
Query: {query}
Candidates:
{candidates_text}
B.4 Paper Information Extraction Prompt